# Robust a posteriori estimation of probit-lognormal seismic fragility curves via sequential design of experiments and constrained reference prior

### v2

@@ 5.2 Estimates and benchmarking metrics
    The 1r1-r square credibility width:
    $\mathcal{W}^{|\mathbf{z}^{k},\mathbf{a}^{k}}=\|q_{1-{r/2}}^{|\mathbf{z}^{k},\mathbf{a}^{k}}-q_{{r/2}}^{|\mathbf{z}^{k},\mathbf{a}^{k}}\|_{L^{2}}^{2}$.
  • 4.
    The square model bias : $\mathcal{M}\mathcal{B}=\|P_{f}^{\text{ref}}-P_{f}^{\text{MLE}}\|_{L^{2}}^{2}$. $P_{f}^{\text{MLE}}$ denotes a probit-lognormal curve whose parameter $\theta$ is obtained by MLE given the full batch of $8\cdot 10^{4}$ simulations.
@@
The model bias $\mathcal{M}\mathcal{B}$ corresponds to the deviations between $P_{f}^{\text{ref}}$ and $P_{f}^{\text{MLE}}$ which can be visualized in Fig. 5. By construction, it is influenced by the distribution of the IM since $P_{f}^{\text{MLE}}$ is estimated by considering all the available samples. We cannot therefore speculate whether this is the absolute minimum bias and its value is only presented for information purposes. We will return to this point in the section devoted to the interpretation of the results.
@@ 5.3.1 Qualitative study
Figure 6: Examples of fragility curves estimation. Top figures: the IM is the PSA; bottom figures: the IM is the PGA. Left figures: the seismic signals are chosen w.r.t. their standard distribution. Right figures: they are chosen w.r.t. our DoE strategy. On each figure: credibility intervals from 2020 (purple) or $80$ (orange) observations, reference fragility curves (green solid line) from Monte Carlo simulations and their $95\%$-confidence intervals (green dashed lines). The red crosses represent the $80$ observations.
@@ 5.3.2 Quantitative study
Figure 7: Probability that a sample of size $k$ yields a degenerate likelihood, as a function of $k$. The considered IM is (left) the PSA and (right) the PGA. The degeneracy probability without DoE (blue dashed line) is compared with the ones with DoE (dotted lines), for different values of $\gamma$. Two extreme values of $\gamma$ are highlighted: $\gamma=0.1$ (purple dotted line) and $\gamma=1.9$ (orange dotted line).
@@
Figure 8: Average error metrics derived from numerous replications of the method, as a function of the size of the observed sample. In dashed blue lines: the average errors using the standard distribution of the IM surrounded by their $95\%$-confidence interval. In dotted lines: the average errors using our DoE strategy with different values of $\gamma$. They are surrounded (in red) by the maximal $95\%$-confidence intervals. Two extreme values are emphasized: $\gamma=0$ (orange) and $\gamma=1.9$ (purple). On the middle figure, the model square bias is plotted as a gray line. The IM is the PSA here.
@@
Up to values of $k$ close to $100$, the DoE method outperforms the standard method. Beyond that, the behavior is different from that observed when the IM is the PSA and an evaluation of the performances of the DoE method is less obvious. On average, however, with the first $40$ data points, the bias value obtained with the PGA is smaller than with the PSA.
@@
Figure 9: As in Figure 8 but here the IM is the PSA.
@@
The practitioner can therefore use these indicators to define a stopping criterion depending on a threshold value. For instance, in our case, a threshold value set at $\mathcal{V}\mathcal{I}_{k}=10^{-3}$ gives approximately $k_{\max}=60$ for the PSA and $k_{\max}=40$ for the PGA. Concerning $\mathcal{V}\mathcal{P}_{k}$, the threshold value can be set at $5\%$. For both the PSA and the PGA, this ensures furthermore that the likelihood degeneracy probability is zero (see Fig. 7). Remark that, even though the number of data required to reach a given stopping criterion is smaller with the PGA than with the PSA, this does not imply that the estimate is less accurate. Indeed, for small data sizes, the degeneracy phenomenon is more likely with the PSA than with the PGA (see Fig. 7). Consequently, the widths of the credibility intervals remain large for large $k$ with the PSA (see Figures 8 and 9). Our approach addresses this phenomenon by reducing the degeneracy even for a small number of observations. In practice, however, this number remains dependent on the IM of interest (see the paragraph “Degeneracy disappearance”).
Figure 10: Average variations of the DoE index $\mathcal{V}\mathcal{I}_{k}$ as a function of the observed sample size $k$, using the PSA (left) and the PGA (right).
Figure 11: Average variations of the median a posteriori fragility curve $\mathcal{V}\mathcal{P}_{k}$ as a function of the observed sample size $k$, using the PSA (left) and the PGA (right).
@@ 5.4 Synthesis
The objective of the methodology proposed in this work is to make the most of the probit-lognormal model. Indeed, it appears to the practitioner as a model that is both pragmatic and relevant for estimating fragility curves with a reduced number of data, judging by the abundant use made of it in the literature. The limitation on the number of data is essential. It comes from the fact that (i) these are expensive in absolute terms, especially with experimental tests, and (ii) the probit-lognormal model being likely to be biased, there is no point in feeding it with a lot of data. To achieve this goal, we proposed a DoE methodology in a Bayesian framework based on the reference prior theory.
Two current IMs —the PSA and the PGA— were considered to study the performance of the method which was applied to an equipment from the nuclear industry. In both cases, with a small number of data —of the order of 80—, the performance of the proposed method was much better than that of the standard one. Performance is best if the method is used with an IM that has a good level of correlation with the structural response of interest.
In order to guide the user, we also proposed two indicators that allow learning to be stopped based on quantitative information. Both criteria appeared sensitive to the change of IM and therefore to the associated potential biases. Let us add, however, that since the degeneracy phenomenon does not make the estimation optimal from the point of view of the size of the credibility interval, it is appropriate, in practice, to continue learning until a non-degenerate sample is obtained (see Definition 1).
@@ 6 Conclusion
Assessing the seismic fragility of structures and components when few binary data are available —i.e., less than 100— is a challenging task. To do this, it is often necessary to introduce some simplifying assumptions. A standard solution found in the literature is to use a parametric model of the fragility curves such as the probit-lognormal model. This is the choice that was made in this work.
@@
In this work, we have favored the Bayesian framework by relying on the reference prior theory in order to define the prior. This framework makes it possible to define a so-called objective prior in the sense that it only depends on the statistical model, namely the probit-lognormal model fed by binary data indicating the state of the structure of interest —i.e., failure or non-failure— for given IM values. The authors showed in a previous work that the solution to this problem is the well known Jeffreys prior. Since it is improper and does not allow to regularize the problem of degeneracy of the likelihood, a proper approximation of it has been proposed in this work. This results in a slightly informed prior that also has the advantage of being completely analytical, which facilitates its numerical implementation. From this prior, authors have proposed a strategy of design of experiments inherited from the reference prior theory, i.e., based on the information theory. Given a large database of synthetic seismic signals, the proposed strategy intends to sequentially select the synthetic signals with which to perform the calculations or the tests, in order to optimally estimate —by minimizing their number— the probit-lognormal estimations of fragility curves.
@@ A.2 Results
Figure 12: For each of the three toy case studies: the distribution of the 100 last selected IM by the method (red histograms) along with the associated reference fragility curve (green) of the toy case study, which is cut at a certain quantile $q$ in each figure. The median of the reference (solid blue) and the values $a^{u}_{1}$, $a^{u}_{2}$ (dashed pink) are emphasized.
@@
Figure 13: Left figure: average square bias of the three toy case studies as a function of the number of observations. Middle figure: index $\mathcal{V}\mathcal{I}_{k}$ for the three toy case studies as a function of the number of observations. Right figure: index $\mathcal{V}\mathcal{P}_{k}$ for the three toy case studies as a function of the number of observations.
@@ A.3 Simple suggestion of a domain where IMs should be theoretically selected
$\hskip-10.0pt\frac{d}{da}\mathbb{E}|P^{\text{ref}}_{f}(a)-P_{f}(a)|=\frac{1}{a\sqrt{2\pi}}\int_{\Theta}\left[\frac{1}{\beta^{\ast}}\exp\left(-\frac{(\log a-\log\alpha^{\ast})^{2}}{2\beta^{\ast 2}}\right)\right.\\ \left.-\frac{1}{\beta}\exp\left(-\frac{(\log a-\log\alpha)^{2}}{2\beta^{2}}\right)\right]p_{\alpha,\beta}(\alpha,\beta)d\alpha d\beta.$ (6)
@@
$p_{\alpha,\beta}(\alpha,\beta)=K(c,d,\tau,\zeta)\left(\frac{1}{\beta^{2}}\right)^{c+1/2}\hskip-10.0pt\exp\left(-\frac{d}{\beta^{2}}\right)\exp\left(-\tau\frac{(\log\alpha-\zeta)^{2}}{2\beta^{2}}\right),$
@@ A.5 Equivalence between toy case studies
$=\Phi\left(\beta^{\ast-1}_{2}\left[\frac{\log c_{2}^{2}-\log c_{2}^{1}}{\log c_{1}^{2}-\log c_{1}^{1}}\left(\log\tilde{a}-\log c_{1}^{1}\right)+\log c_{1}^{1}-\log\alpha^{\ast}_{2}\right]\right),$

### v3

@@ 5.2.2 Benchmarking metrics
    The 1r1-r square credibility width: $\mathcal{W}^{|\mathbf{z}^{k},\mathbf{a}^{k}}=\|q_{1-{r/2}}^{|\mathbf{z}^{k},\mathbf{a}^{k}}-q_{{r/2}}^{|\mathbf{z}^{k},\mathbf{a}^{k}}\|_{L^{2}}^{2}$.
@@
5.2.3 Model bias
In practice, the probit-lognormal model is susceptible to be biased. So, one of the metrics for evaluating its performance is to assess the square bias to the median, $\mathcal{B}^{|\mathbf{z}^{k},\mathbf{a}^{k}}$, defined in the previous section. To facilitate the interpretation of the values provided by this metric, we propose estimating the square model bias, $\mathcal{M}\mathcal{B}$, which corresponds to the deviations between $P_{f}^{\text{ref}}$ and $P_{f}^{\text{MLE}}$. It can be visualized in Fig. 5 and is defined by:
$\mathcal{M}\mathcal{B}=\|P_{f}^{\text{ref}}-P_{f}^{\text{MLE}}\|_{L^{2}}^{2}.$ (6)
$P_{f}^{\text{MLE}}$ denotes a probit-lognormal curve whose parameter $\theta$ is obtained by MLE given the full batch of $8\cdot 10^{4}$ simulations. Obviously, this curve would not be known in a real situation since its knowledge is equivalent to the knowledge of $P_{f}^{\text{ref}}$. It serves the assessment of our method: $P_{f}^{\text{MLE}}$ is expected to approach the “best possible” curve among probit-lognormal ones for this case study. By construction, the model bias is influenced by the distribution of the IM since $P_{f}^{\text{MLE}}$ is estimated by considering all the available samples. We cannot therefore speculate whether this is the absolute minimum bias and its value is only presented for information purposes. We return to this point in the section devoted to the interpretation of the results.
@@ 5.3 Numerical results
Top figures: the IM is the PSA; bottom figures: the IM is the PGA. Left figures: the seismic signals are chosen w.r.t. their standard distribution. Right figures: they are chosen w.r.t. our DoE strategy. On each figure: credibility intervals from $20$ (purple) or $80$ (orange) observations, reference fragility curves (green solid line) from Monte Carlo simulations and their $95\%$-confidence intervals (green dashed lines). The red crosses represent the $80$ observations.
Figure 6: Examples of fragility curves estimation.
@@ 5.3.2 Quantitative study
Left: the considered IM is the PSA. Right: it is the PGA. The degeneracy probability without DoE (blue dashed line) is compared with the ones with DoE (dotted lines), for different values of $\gamma$. Two extreme values of $\gamma$ are highlighted: $\gamma=0.1$ (purple dotted line) and $\gamma=1.9$ (orange dotted line).
Figure 7: Probability that a sample of size $k$ yields a degenerate likelihood, as a function of $k$.
@@
As shown in Fig. 9, up to values of $k$ close to $100$, the DoE method outperforms the standard method. Beyond that, the behavior is different from that observed when the IM is the PSA and an evaluation of the performances of the DoE method is less obvious. On average, however, with the first $40$ data points, the bias value obtained with the PGA is smaller than with the PSA.
@@
In dashed blue lines: the average errors using the standard distribution of the IM surrounded by their $95\%$-confidence interval. In dotted lines: the average errors using our DoE strategy with different values of $\gamma$. They are surrounded (in red) by the maximal $95\%$-confidence intervals. Two extreme values are emphasized: $\gamma=0$ (orange) and $\gamma=1.9$ (purple). On the middle figure, the model square bias is plotted as a gray line. The IM is the PSA here.
Figure 8: Average error metrics derived from numerous replications of the method, as a function of the size of the observed sample.
Figure 9: As in Figure 8 but here the IM is the PGA.
@@
The practitioner can therefore use these indicators to define a stopping criterion depending on a threshold value. For instance, in our case, a threshold value set at $\mathcal{V}\mathcal{I}_{k}=10^{-3}$ gives approximately $k_{\max}=60$ for the PSA and $k_{\max}=40$ for the PGA. Concerning $\mathcal{V}\mathcal{P}_{k}$, the threshold value can be set, for instance, at $5\%$: beyond this point, the index indicates that the estimated fragility curve exhibits minimal evolution when additional data are generated. For both the PSA and the PGA, this also guarentees that the likelihood degeneracy probability is zero (see Fig. 7). These thresholds reflect our interpretation of the present case study and are not supported by existing guidance in the literature; practitioners should therefore select appropriate thresholds based on their objectives and the specific characteristics of the system they study. Remark that, even though the number of data required to reach a given stopping criterion is smaller with the PGA than with the PSA, this does not imply that the estimate is less accurate. Indeed, for small data sizes, the degeneracy phenomenon is more likely with the PSA than with the PGA (see Fig. 7). Consequently, the widths of the credibility intervals remain large for large $k$ with the PSA (see Figures 8 and 9). Our approach addresses this phenomenon by reducing the degeneracy even for a small number of observations. In practice, however, this number remains dependent on the IM of interest (see the paragraph “Degeneracy disappearance”).
Left: the considered IM is the PSA. Right: the considered IM is the PGA.
Figure 10: Average variations of the DoE index $\mathcal{V}\mathcal{I}_{k}$ as a function of the observed sample size $k$
Left: the considered IM is the PSA. Right: the considered IM is the PGA.
Figure 11: Average variations of the median a posteriori fragility curve $\mathcal{V}\mathcal{P}_{k}$ as a function of the observed sample size $k$.
@@ 5.4 Synthesis
Judging by its abundant use in the literature, the probit-lognormal model is nowadays considered to the practitioner as a model that is both pragmatic and relevant for estimating fragility curves. Its use with a limited amount of data is nevertheless a crucial point. Besides the facts that data is generally expensive —especially for experimental tests— and that the probit-lognormal model is likely to be biased in practice, there is no point in feeding it with a large amount of data. The objective of the methodology proposed in this work is then to make the most of the probit-lognormal model by obtaining the most accurate estimate possible with the minimum amount of data. To this end, we proposed a DoE methodology in a Bayesian framework based on the reference prior theory.
In practice, the choice of the best IM to consider for expressing fragility curves is an open subject. Although some criteria exist in the literature [6, 8, 78], there is no unique solution, as the choice ideally depends on the structure being studied. In practice, this choice may therefore be dictated by practical considerations. For example, at the scale of a large industrial facility, such as a nuclear power plant, the PGA may be chosen as IM for all relevant equipment to facilitate risk assessment studies, even if it is not optimal.
Although in A it is demonstrated, through a toy case study, that in theory the performance of the DoE method is not affected by the choice of the IM, this is not the case in practice. What primarily affects its performance is the bias of the model, which can differ from one IM to another. So, in this work, two current IMs —the PSA and the PGA— were considered to assess the performance of the method which was applied to an equipment from the nuclear industry. Thus, it is shown that overall performance is best if the method is used with an IM that has a good level of correlation with the structural response of interest, such as the PSA in our setting. Nevertheless, in both cases, with a small number of data —of the order of 80—, the performance of the proposed method is much better than that of the standard one. In our study, when the PGA is chosen as the IM, the method’s performance degrades and is worse than with the standard method beyond 100 data points. This performance degradation is explained by the fact that the bias is more pronounced with the PGA in one of the two main domains where the DoE method selects data.
These results reinforce the idea of using the probit-lognormal model only with small datasets, due to the inherent model bias that can be expected in practice. In order to guide the user, we also proposed two indicators that allow learning to be stopped based on quantitative information. Both criteria appeared sensitive to the change of IM and therefore to the associated potential biases. Let us add, however, that since the degeneracy phenomenon does not make the estimation optimal from the point of view of the size of the credibility interval, it is first of all appropriate, in practice, to continue learning until a non-degenerate sample is obtained (see Definition 1).
@@ 6 Conclusion
Assessing the seismic fragility of structures and components when few binary data are available —i.e., less than 100— is a challenging task. To do this, it is often necessary to introduce some simplifying assumptions. A standard solution is to use a parametric model of the fragility curves such as the probit-lognormal model. This is the choice that was made in this work.
@@
In this work, we have favored the Bayesian framework by relying on the reference prior theory in order to define the prior. This framework makes it possible to define a so-called objective prior in the sense that it only depends on the statistical model, namely the probit-lognormal model fed by binary data indicating the state of the structure of interest —i.e., failure or non-failure— for given IM values. The authors showed in a previous work that the solution to this problem is the well known Jeffreys prior. Since it is improper and does not allow to regularize the problem of degeneracy of the likelihood, a proper approximation of it has been proposed in this work. This results in a slightly informed prior that also has the advantage of being completely analytical, which facilitates its numerical implementation. The proposed analytical expression is obtained by assuming a lognormal distribution for the IM. This assumption is consistent with the real seismic signals considered in this study, which correspond to a specific seismic scenario, but the literature shows that it is suitable for more general applications (near-source ground motions, artificial seismic signals generated from GMPE, etc.). Thus, thanks to its simple analytical formulation and its wide field of application, this prior can prove very useful in engineering practice using the Bayesian framework. From this prior, authors have proposed a strategy of design of experiments inherited from the reference prior theory, i.e., based on the information theory. Given a large database of synthetic seismic signals, the proposed strategy intends to sequentially select the synthetic signals with which to perform the calculations or the tests, in order to optimally estimate —by minimizing their number— the probit-lognormal estimations of fragility curves.
@@ A.2 Results
For each of the three toy case studies: the distribution of the 100 last selected IM by the method (red histograms) along with the associated reference fragility curve (green) of the toy case study, which is cut at a certain quantile $q$ in each figure. The median of the reference (solid blue) and the values $a^{u}_{1}$, $a^{u}_{2}$ (dashed pink) are emphasized.
Figure 12: Stationary distributions of the selected IM by the method.
@@
Figure 13: Average values of $\mathcal{B}^{|\mathbf{z}^{k},\mathbf{a}^{k}}$, $\mathcal{V}\mathcal{I}_{k}$, $\mathcal{V}\mathcal{P}_{k}$ as functions of the number of observation $k$ for the three toy case studies.
@@ A.3 Simple suggestion of a domain where IMs should be theoretically selected
$\hskip-5.0pt\frac{d}{da}\mathbb{E}|P^{\text{ref}}_{f}(a)-P_{f}(a)|=\frac{1}{a\sqrt{2\pi}}\int_{\Theta}\left[\frac{1}{\beta^{\ast}}\exp\left(-\frac{(\log a-\log\alpha^{\ast})^{2}}{2\beta^{\ast 2}}\right)\right.\\ \left.-\frac{1}{\beta}\exp\left(-\frac{(\log a-\log\alpha)^{2}}{2\beta^{2}}\right)\right]p_{\alpha,\beta}(\alpha,\beta)d\alpha d\beta.$ (7)
@@
$p_{\alpha,\beta}(\alpha,\beta)=K(c,d,\tau,\zeta)\left(\frac{1}{\beta^{2}}\right)^{c+1/2}\hskip-8.99994pt\exp\left(-\frac{d}{\beta^{2}}\right)\exp\left(-\tau\frac{(\log\alpha-\zeta)^{2}}{2\beta^{2}}\right),$
@@ A.5 Equivalence between toy case studies
$=\Phi\left(\beta^{\ast-1}_{2}\left[\frac{\log c_{2}^{2}-\log c_{2}^{1}}{\log c_{1}^{2}-\log c_{1}^{1}}\left(\log\tilde{a}-\log c_{1}^{1}\right)+\log c_{1}^{1}-\log\alpha^{\ast}_{2}\right]\right)$